OpenStudy (anonymous):

Σ n = 1(start value) ∞(end value) 6 ( – 3 5 ) n(sequence) a. – 9/4 b. 60/7 c. –9 d. 15 <---i choose this answer

OpenStudy (amistre64):

-1 + x - x^2 + x^3 - x^4 + .... this is a "tricky" series. -1 + x ------------------------ 1+x ) -1 (-1-x) ------ x (x+x^2) -------- -x^2 but this seems to be it: -1/(1+x)

OpenStudy (amistre64):

let x = 3/5

OpenStudy (amistre64):

\[\sum_1^{inf}6(-\frac{3}{5})^n\] \[6(3)\sum_1^{inf}(-\frac{1}{5})^n\] \[6(3)~\frac 15(-\frac{1}{1+1/5})\] \[6(3)~\frac 15(-\frac{1}{6/5})\] \[\cancel 6(3)~\frac 1{\cancel 5}(-\frac{\cancel 5}{\cancel 6})=-3\] gonna have to rethink this :)

OpenStudy (amistre64):

the wolf says it aint d

OpenStudy (amistre64):

ahh, i took out the 3 trying to be smart lol

OpenStudy (anonymous):

:P

OpenStudy (amistre64):

\[6~\frac 35(-\frac{1}{1+3/5})\] \[-6~\frac 35(\frac{1}{8/5})\] \[-6~\frac 35(\frac{5}{8})\] \[-3~\frac 35(\frac{5}{4})\] thats better

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sum%28n%3D1..inf%29+%286%28-3%2F5%29%5En%29 now the wolf agrees :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

how would i solve this question? f(x) = 3x^2 – 5x; x_0 = 1 a. 4, 11, 19, 28, 38, 49 b. 4, 10, 16, 22, 28, 34 c. 4, 12, 21, 31, 42, 54 d. 4, 10, 18, 27, 37, 48

OpenStudy (amistre64):

it seems to be just letting x=1,2,3,4,5 and finding the results

OpenStudy (amistre64):

or, it seems that there might be some content of the question that is missing

OpenStudy (anonymous):

oops sorry the correct multiple choices are: a. –2, 7, 142 b.–1, –2, –11 c. 2, 5, 63 d. –2, –11, –38

OpenStudy (amistre64):

well, when x=1, we get f=-2 but nothing else seems to fit.

OpenStudy (amistre64):

f(x) = 3x^2 – 5x; x_0 = 1 this suggests that there is some iterative process going on but it is lacking any clear instructions to put it into context

OpenStudy (anonymous):

its says to find the first three iterates of the function for the given initial value

OpenStudy (amistre64):

well, -2,2,12 doesnt seem to be an option, and it doesnt seem clear yet that it is asking for a Newton method to find zeros ...

OpenStudy (amistre64):

-2, 22 ,... would be the next reasonable approach as is, but even that is not an option

OpenStudy (anonymous):

yeah i kept getting -2,22

OpenStudy (amistre64):

f(x_0), f(f(x_0)), f(f(f(x_0))) is how an iterative process would work f(1), f(f(1)), f(f(f(1))) -2, f(-2), f(f(-2)) -2, 22, f(22)